OpenStudy (anonymous):

The sum of three consecutive integers is -165, What is the closest to zero?

OpenStudy (aum):

Let the three consecutive numbers be: (n-1), n, (n+1) Sum = (n-1) + n + (n+1) = 3n = -165 solve for n. Find the 3 numbers and the one closest to zero.